OpenStudy (anonymous):

Find the points on the curve below at which the tangent is horizontal. Use n as an arbitrary integer. (Enter your answers as a comma-separated list.) y = sinx/(2+cosx)

OpenStudy (anonymous):

take the derivative, set it equal to zero and solve. how far have you gotten?

OpenStudy (anonymous):

just got the derivative which is \[y'= (2\cos(x)+1)/(\cos(x)+2)^2\]

OpenStudy (anonymous):

looks good to me. so set \[2\cos(x)+1=0\] \[\cos(x)=-\frac{1}{2}\] and so we get an infinite number of possibilities for x

OpenStudy (anonymous):

lets see if we can find a formula for them. \[\cos(\frac{2\pi}{3})=-\frac{1}{2}\] and so one answer would be \[\frac{2\pi}{3}+2k\pi, k \in \mathbb Z\]

OpenStudy (anonymous):

another answer is \[\cos(\frac{4\pi}{3})=-\frac{1}{2}\] so we could also write \[x=\frac{4\pi}{3}+2k\pi, k \in \mathbb Z\]

OpenStudy (anonymous):

i don't see a way to combine this into one formula, but maybe that is because i am not smart enough.